OpenStudy (anonymous):

What is the solution to the rational equation 4/3 + 2/x-5 = -6/3x-15

OpenStudy (pfenn1):

\[\frac43+\frac{2}{x+5}=\frac{-6}{3x-15}\]Find the LCD. If you need some review on how to do this, check this very helpful website http://www.helpwithfractions.com/least-common-denominator.html